# Neil A. Chriss

### v2

@@ 2.6 Future directions
Finally, we give a brief discussion of various directions one might consider. This paper’s primary aim is to shed light on strategy decision making while trading in competition and we believe a game-theoretic framework is the right starting point. However, it is obvious that this framework has many shortcomings that provide a multitude of potential research directions. The good news is that the historical development of game theory itself provides a useful template for how one might evolve this research.
@@
Each item in this list marks a potential direction of research by going beyond our framework in a direction that tracks the development of game theory since Nash. For example, traders repeatedly trade in competition, so to what extent would modeling this as a repeated game impact equilibria? Also, in reality traders do not "make their moves" simultaneously, but rather different traders not only start their position-building at different times, but at times they go so far as to announce their ideas in conferences, public interview and private "idea dinners" with other investors. In addition, modeling this as a perfect information game is decidedly not a perfect information game. Traders do not know what the other traders’ strategies are or what moves have occurred up to a given time.
@@
In addition to the game theoretic limitations, there is also the question which market impact models should one use to inform trading in competition? At this point it may seem like the only right answer is to use models that most accurately predict the impact of trading on prices, however a counterpoint to this is that since we are not, in fact, interested in predicting market impact or implementation cost itself, it is not wholly obvious that this is the right answer. One reason for this is because the details of market impact, especially alpha-decay will vary over time and must be estimated each time. Therefore, it is worth considering robust models that prioritize minimizing the cost of errors in parameters over the specificity of the models. Despite this, it is still worth examining the how other models might change the shape of strategies in equilibrium and importantly whether or not the nature of equilibrium itself is dependent on the market impact models used.
@@ 2.7 Organization of this paper
Here we give a brief overview of the rest of this paper. In Section 3 we examine what happens when all traders’ strategies are consolidated into a single strategy. We do this so that later we can examine how the associated costs of a market minimized strategy compare with those in equilibrium. Next, in Section 4 we study the market strategy, the aggregation of all strategies treated as if it were a strategy, and we study its properties at equilibrium. In Section 5 we prove that equilibrium strategies exist and in Proposition 5.1 we provide closed form solutions for those strategies. This section also derives formulas for implementation costs in equilibrium. Finally, in Section 6 we examine trade centralization and study how and to what extent it can improve implementation costs for a firm that employs multiple traders. Finally, we relegate most of the proofs of purely mathematical propositions to a set of appendices at the end of this paper.
3 Whole market cost minimization
@@
We examine this here so that later we can analyze the price of anarchy referred to in the introduction. While the result is immediate if we simply treat the market strategy as a single strategy, we show here its precise derivation as the sum of the implementation costs of each strategy. Let $C$ be the total cost of trading (function) equal to the sum of the individual costs $C_{i,\lambda_{i}}$. In that each $C_{i,\lambda_{i}}$ is equal to $(\dot{m}+\kappa m)\lambda_{i}\dot{a}_{i}$ we immediately see
@@
This relation depends on the fact that the market impact is linear. Namely for two general strategies $x$ and $y$ the cost of trading $x$ and $y$ simultaneously is equal to the sum costs of trading each. With that said we apply the Euler-Lagrange equation to find the strategy that minimizes implementation cost.
@@
Proposition 3.1.
If the central trading desk wishes to minimize costs it will trade the risk-neutral strategy $m(t)=t$ and the implementation cost (the aggregate cost) will be the integral of $(\dot{m}+\kappa m)\dot{m}$ and since $\dot{m}=1$ we have
Aggregate $\text{Aggregate cost}=\int_{0}^{1}1+\kappa t\,\mathrm{dt}=1+\frac{\kappa}{2}$ (3.4)
As this represents the implementation cost to the central trading desk, if trader $i$’s target quantity is $\lambda_{i}$ where $\lambda_{i}$ is the fraction of the total quantity being acquired, then trader $i$’s share of the costs will be $\lambda_{i}(1+\frac{\kappa}{2})$. In itself this is interesting but later in Corollary 4.2 be able to give a precise comparison of this cost to the total implementation cost in the case where each trader acts independently and will see that the centralization represents a potentially significant savings. $\blacksquare$
@@ 4 The market strategy
In this section we return to the general situation of an $n$-trader non-cooperative game. That is, we assume each of the $n$-traders is merely trying to minimize costs relative to the loss function $C_{i,\lambda_{i}}(t)$. In this case we are able to give a complete description of the equilibrium.
@@ 4.1 The market strategy is eager
Let $\lambda_{1},\dots,\lambda_{n}$ be the scaled target quantities of $n$ traders in competition to buy the same stock. If $a_{\lambda_{1}},\dots,a_{\lambda_{n}}$ are a Nash equilibrium then the market strategy $m(t)=\sum_{i=1}^{n}a_{\lambda_{i}}(t)$ is the strategy:
@@
An immediate corollary of this is that the total implementation cost depends only on $\kappa$ and the number of traders $n$. To see this we only need to note that the total cost is the sum of the individual cost functions. Recalling that aggregate cost is the sum over all $i$ of $C_{i,\lambda_{i}}(t)=\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}\dot{a}_{i}(t)$ and writing
Aggregate $\begin{split}\text{Aggregate cost}&=\sum_{i=1}^{n}\int_{0}^{1}C_{i,\lambda_{i}}(t)\,\mathrm{dt}\\ &=\sum_{i=1}^{n}\int_{0}^{1}\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}\dot{a}_{i}(t)\,\mathrm{dt}=\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\dot{m}(t)\,\mathrm{dt}\end{split}$ (4.2)
@@
$\lim_{n\to\infty}\text{Aggregate cost}=\frac{\kappa}{e^{\kappa}-1}+\kappa=\frac{\kappa e^{\kappa}}{e^{\kappa}-1}$ (4.4)
@@
This allows us to immediately compute the price of anarchy which is the limiting value as $n\to\infty$ of the ratio of the aggregate cost to the market minimizing cost (see Section 3 $(1+\frac{\kappa}{2})^{-1}$. Figure 3 display the relationship between the price of anarchy, $\kappa$ and the number of traders.
@@
Thus we conclude that the price of anarchy is bounded above by $2$ and this bound is tight. $\blacksquare$
Figure 3: The price of anarchy as it relates to the level alpha-decay and the number of traders. Roughly speaking, the greater the price of anarchy the more the market as a whole benefits from centralization. Equally true is that the greater the $\kappa$, that is the more alpha-decay there is, the more the marfet benefits from centralization.
@@ 5 The equilibrium strategies
In this section compute the equilibrium strategies and explore some of their distinctive properties including the cost of trading in an equilibrium.
@@
Suppose there are $n$ traders labeled $i=1,\dots,n$ trading in competition and that trader $i$’s target quantity is the fraction $10<\lambda_{i}<$ 1 of the total target quantity. Then if each trader $i$ trades the strategy:
@@
and further the resultant market strategy $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ derived by summing the $a_{i}$ in Eq. 5.1 again yields the eager strategy given in Proposition 4.1. In next Section 5.2 we show that for the case that $\lambda_{i}=1/n$ for all $i$ each trader traders a $1/n$-scaled replica of this strategy. Also in proposition 5.2 we will compute the cost of trading each equilibrium strategy.
@@ 5.3 The cost of trading in equilibrium
$\text{Cost}(\lambda_{i}a_{i},m)=\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}$ (5.4)
We give the proof of this in Appendix A.4 but note a few facts here. First, the proof exploits the relationship between $\ddot{a}_{i}-\kappa\dot{a}_{i}$ and $\dot{m}$, without which appears to be very difficult. Second, summing $\text{Cost}(\lambda_{i}a_{i},m)$ over all strategies $i=1,\dots,n$ yields the total implementation cost, or simply aggregate cost. We show in Appendix A.4 that this sum is equal to:
Aggregate $\text{Aggregate cost}=\sum_{i=1}^{n}\text{Cost}(\lambda_{i}a_{i},m)=\frac{\alpha}{e^{\alpha}-1}+\frac{\kappa n}{n+1}$ (5.5)
and while it is not immediately obvious, this is equivalent to the formula for the total implementation cost in Corollary 4.2. We also note that since as $n\to\infty$, $\alpha\to\kappa$, total implementation cost tends to $\frac{\kappa}{e^{\kappa}-1}+\kappa=\frac{1}{1-e^{-\kappa}}$, which is the same as 4.4. $\blacksquare$
@@
5.4 Each trader’s implementation as a share of the total
We have already seen that the total implementation costs over all traders does not depend on the distribution of target quantities, despite the fairly large differences between the shapes of each strategy. Therefore it’s reasonable to ask whether each traders share of total cost is simply the their expected share of the total cost. For example, if a trader is trading 25% of the total quantity it might be the case that their share of the total cost is simply 25% of the total. We answer this firmly in the negative in Proposition 5.3.
@@ 6 Centralized trading
$\underbrace{\lambda_{1},\dots,\lambda_{n_{1}}}_{\text{Firm, $\lambda_{n_{1}}$}}\quad\text{and}\quad\underbrace{\lambda_{n_{1}+1},\dots,\lambda_{n_{2}}}_{\text{Non-firm, $\lambda_{n_{2}}$}}$ (6.1)
Each trader within the firm has a target quantity and associated parent order associated they would like to execute (see Section 2.2. If the firm funnels all the parent orders through an internally operated trading desk, the central trading desk, then it is said to be centralizing its trades. A prominent reason for trade centralization is to reduce implementation costs. The thinking behind this is that if each trader trades independently then they will operate to minimize their own implementation costs.
@@
Given this setup there are two possible approaches the firm can take. The first case is that they can allow each trader to trade independently (the non-cooperative approach) and the second is the firm can centralize all the trades (the cooperative approach). We summarize as in the following table:
@@
Non-cooperative $n$ $n_{1}$ $\lambda_{n_{1}}$ $n_{2}$ $\lambda_{n_{2}}$
Cooperative $n_{2}+1$ 1 $\lambda_{n_{1}}$ $n_{2}$ $\lambda_{n_{2}}$
Table 1: Firm and non-firm trading setup as it pertains to how the parameters of competitive trading change when moving from cooperative to non-cooperative trading.
@@ 6.1 Naive centralization
When a firm naively centralizes it implicitly takes the view that the best way to minimize implementation costs is to trade the centralized order flow as a single monolithic strategy designed to minimize its total implementation cost. In this section, then, we study what this strategy looks like in equilibrium and compare its implementation cost to that of the de-centralized strategies (also in equilibrium). The parameters of the trading game immediately become those of the cooperative row of Table 1. Given this, there are two instances of the firm strategy: one when the traders trade independently and the other where their trades are centralized. We refer to these as the non-cooperative and cooperative strategies respectively.
@@
$\lambda_{n_{1}}a_{n_{1}}$ $=\lambda_{n_{1}}(B_{n_{1}}(e^{\kappa t}-1)+D_{n_{1}}(1-e^{-\hat{\alpha}t})),\qquad\text{\em Cooperative firm strategy}$ (6.2b)
@@
and $\alpha=\kappa\frac{n-1}{n+2}$, $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+1}$ and $n_{1}+n_{2}=n$. Next there are the corresponding strategies for the non-firm traders:
@@
In Figure 7 we display comparisons of the firm and non-firm strategies in the cooperative versus the non-cooperative cases. In each plot green lines represent equilibrium aggregated strategies when firm traders centralize their trades (the cooperative case) and blue lines represent the same but for the non-cooperative (when traders do not centralize). Solid lines represent the firm’s aggregated strategy in equilibrium and dashed lines represent the non-firm traders’ aggregated strategy. We can think of a pair of lines of the same color (one solid and one dashed) as a strategy and a counter-strategy. We now discuss the meaning of this.
@@
Figure 8: Another example of firm versus non-firm aggregated strategies in cooperative (centralized) versus non-cooperative situations, see Figure 7.
Figures 7 and 8 suggest a surprising hypothesis. Firms that consolidate trades may not be gaining by doing so. We examine this further next.

### v3

@@ 2.7 Future directions
To conclude this section, we give a brief discussion of various directions one might consider. This paper’s primary aim is to shed light on strategy decision making while trading in competition and we believe a game-theoretic framework is the right starting point. However, it is obvious that this framework has many shortcomings that provide a multitude of potential research directions. The good news is that the historical development of game theory itself provides a useful template for how one might evolve this research.
@@
Each item in this list marks a potential direction for new research by going beyond the framework in this paper. In so doing, this takes the research in a direction that tracks the development of game theory since Nash. For example, traders repeatedly trade in competition, so to what extent would modeling this as a repeated game impact equilibrium strategies? Also, in reality traders do not "make their moves" simultaneously, but rather different traders not only start their position-building at different times, but at times they go so far as to announce their ideas in conferences, public interview and private "idea dinners" with other investors. In addition, modeling this as a perfect information game is decidedly not a perfect information game. Traders do not know what the other traders’ strategies are or what moves have occurred up to a given time.
@@
In addition to new game theoretic directions, there is also the question which market impact models should one use to inform trading in competition? After all, this work is not exclusively of academic interest. We should be asking ourselves how different market impact models impact equilibrium strategies and which ones would yield the best results in practice. At this juncture it may seem that the only correct answer is to use models that most accurately predict the impact of trading on prices, however a counterpoint to this is that since we are not, in fact, interested in predicting market impact or implementation cost itself, it is not wholly obvious that this is the right answer. One reason for this is because the details of market impact, especially alpha-decay will vary over time and must be estimated each time. Therefore, it is worth considering robust models that prioritize minimizing the cost of errors in parameters over the specificity of the models. Despite this, it is still worth examining the how other models might change the shape of strategies in equilibrium and importantly whether or not the nature of equilibrium itself is dependent on the market impact models used.
@@ 2.8 Organization of this paper
The rest of this paper is organized as follows. In Section 3 we examine what happens when all traders’ strategies, market-wide, are consolidated into a single strategy. We do this to examine how the associated costs of a market-wide strategy compare with those in the equilibrium where every trader acts independently to minimize their costs. After that, in Section 4 we study the market strategy, the aggregation of all traders’ strategies treated as if it were a single strategy. We study the key properties of the market strategy in equilibrium and show that the market strategy is always an eager strategy. In Section 5 we prove that equilibrium strategies exist and in Proposition 5.1 we provide closed form solutions for those strategies. This section also derives formulas for implementation costs in equilibrium. Finally, in Section 6 we examine trade centralization and study how and to what extent it can improve implementation costs for a firm that employs multiple traders. Finally, we relegate most of the proofs of purely mathematical propositions to a set of appendices at the end of this paper.
3 Market-wide cost minimization
@@
We examine this here so that later we can analyze the price of anarchy referred to in the introduction. While the result is immediate if we simply treat the market strategy as a single strategy, we show here its precise derivation as the sum of the implementation costs of each strategy. Let $C$ be the total cost of trading (function) equal to the sum of the individual costs $C_{i}$. In that each $C_{i}$ is equal to $(\dot{m}+\kappa m)\lambda_{i}\dot{a}_{i}$ we immediately see
@@
This relation depends on the fact that the market impact is additive. More precisely, the relationship replies on the fact that we are modeling the market impact of collective trading to be the super-position of the market impact of individual trading. Namely for two general strategies $x$ and $y$ the cost of trading $x$ and $y$ simultaneously is equal to the sum costs of trading each. With that said we apply the Euler-Lagrange equation to find the strategy that minimizes implementation cost.
@@
Proposition 3.1 (Market-wide cost minimization).
If a "market-wide" central trading desk minimizes aggregate costs, it will trade the risk-neutral strategy $m(t)=t$ and the implementation cost (the aggregate cost) will be the integral of $(\dot{m}+\kappa m)\dot{m}$ and since $\dot{m}=1$ we have
[…]
As this represents the implementation cost to a kind-of benevolent market-wide central trading desk, it also says that each trader’s cost is a proportion of the aggregate cost according to their share of the target quantity. That is, if trader $i$’s target quantity is $\lambda_{i}$ (where $\lambda_{i}$ is the fraction of the total quantity being acquired), then trader $i$’s costs is $\lambda_{i}(1+\frac{\kappa}{2})$. In itself this is interesting but later in Corollary 4.2 be able to give a precise comparison of this cost to the total implementation cost in the case where each trader acts independently and will see that the centralization represents a potentially significant savings. $\blacksquare$
@@ 4 The market strategy
In this section we return to the general situation of an $n$-trader non-cooperative game. That is, we assume each of the $n$-traders is merely trying to minimize costs relative to the loss function $C_{i}(t)$. In this case we are able to give a complete description of the equilibrium.
@@ 4.1 The market strategy is eager
Let $\lambda_{1},\dots,\lambda_{n}$ be the scaled target quantities of $n\geq 2$ traders in competition to buy the same stock. If $a_{\lambda_{1}},\dots,a_{\lambda_{n}}$ are a Nash equilibrium then the market strategy $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ is the strategy:
@@
This implies an almost immediate corollary. We see that Eq. 4.1 only involves the market impact parameter $\kappa$ and the number of traders $n$. This means that the total cost of trading the market strategy (see below, Eq. 4.2 can only depend on those parameters. Therefore, the aggregate cost of trading over all traders, that is, the sum of the cost to each trader, does not depend on the mix of target quantities among traders.
We can see this explicitly by examining the sum of the individual cost functions. Recalling that aggregate cost is the sum over all $i$ of $C_{i}(t)=\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}\dot{a}_{i}(t)$ and writing
Aggregate cost (no $\begin{split}\text{Aggregate cost (no centralization)}&=\sum_{i=1}^{n}\int_{0}^{1}C_{i}(t)\,\mathrm{dt}\\ &=\sum_{i=1}^{n}\int_{0}^{1}\big{(}\dot{m}(t)+\kappa\,m(t)\big{)}\lambda_{i}\dot{a}_{i}(t)\,\mathrm{dt}\\ &=\int_{0}^{1}\left(\dot{m}(t)+\kappa\,m(t)\right)\dot{m}(t)\,\mathrm{dt}\end{split}$ (4.2)
@@
$\begin{split}\lim_{n\to\infty}\,\text{Aggregate cost}&=\frac{\kappa}{e^{\kappa}-1}+\kappa\\ &=\frac{\kappa+\kappa(e^{\kappa}-1)}{e^{\kappa}-1}\\ &=\frac{\kappa e^{\kappa}}{e^{\kappa}-1}\\ &=\frac{\kappa}{1-e^{-\kappa}}\end{split}$ (4.4)
@@
This allows us to immediately compute the price of anarchy, the limiting value as $n\to\infty$ of the ratio of the aggregate cost to the market minimizing cost (see Section 3). Figure LABEL:fig:anarchy display the relationship between the price of anarchy, $\kappa$ and the number of traders.
@@ 5 The equilibrium strategies
In this section we compute general equilibrium strategies as in Definition 2.1 and explore some of their distinctive properties including deriving closed-form solutions for the cost of trading in an equilibrium.
@@
Suppose there are $n$ traders labeled $i=1,\dots,n$ trading in competition and that trader $i$’s fractional target quantity is $\lambda_{i}$. Then the strategies $a_{1},\dots a_{n}$ defined directly below represent an equilibrium with respect to the cost functions $C_{1},\dots,C_{n}$ defined in Eq. 2.5:
@@
and further the resultant market strategy $m(t)=\sum_{i=1}^{n}\lambda_{i}a_{i}(t)$ derived by summing the $a_{i}$ in Eq. 5.1 again yields the eager strategy given in Proposition 4.1. In particular, this means that these strategies simultaneously minimize $\int_{0}^{1}C_{i}(t)\,\mathrm{dt}$.
In next Section 5.2 we show that for the case that $\lambda_{i}=1/n$ for all $i$ each trader traders a $1/n$-scaled replica of this strategy. Also in proposition 5.2 we will compute the cost of trading each equilibrium strategy.
@@ 5.1 Visualizing equilibria
Figure 3: Each plot shows the same mix of fractional target quantities but for different levels of $\kappa$. All plots show scaled final quantities, for example, while all plots meet at one unit at time 1, in reality strategy one would acquire 20% of the total. Below each plot the implementation cost to each trader is displayed along with the percentage of the total cost it represents. These are calculated using the formula for total cost given in Section A.4. In each panel, the value of $\lambda_{1}$ (trader 1’s percentage of total target quantity) is set to 0.2 and this corresponding strategy is in blue.
@@ 5.3 The cost of trading in equilibrium
$\text{Cost}(\lambda_{i}a_{i})=\kappa\frac{\lambda_{i}n-1}{n(1-e^{-\kappa})}+\alpha\frac{1}{n(e^{\alpha}-1)}+\frac{\kappa}{n+1}$ (5.4)
We give a proof of this in Appendix A.4 but note a few facts here. First, the proof exploits the relationship between $\ddot{a}_{i}-\kappa\dot{a}_{i}$ and $\dot{m}$, without which appears to be very difficult. Second, summing $\text{Cost}(\lambda_{i}a_{i},m)$ over all strategies $i=1,\dots,n$ yields the total implementation cost, or simply aggregate cost. We show in Appendix A.4 that this sum is equal to:
Aggregate $\text{Aggregate cost}=\sum_{i=1}^{n}\text{Cost}(\lambda_{i}a_{i})=\frac{\alpha}{e^{\alpha}-1}+\frac{\kappa n}{n+1}$ (5.5)
and while it is not immediately obvious, this is equivalent to the formula for the total implementation cost in Corollary 4.2.
We also note that since as $n\to\infty$, $\alpha\to\kappa$, and therefore the aggregate cost expression Eq. 5.5 tends to $\frac{\kappa}{e^{\kappa}-1}+\kappa=\frac{\kappa}{1-e^{-\kappa}}$, which is the same as 4.4. $\blacksquare$
@@
5.4 Each trader’s implementation cost as a share of the total cost
We have already seen that the total implementation costs over all traders does not depend on the distribution of target quantities, a somewhat surprising fact given the fairly large range of possible strategy shapes. Nevertheless it’s reasonable to ask whether each traders share of total cost is simply their expected share of the total cost. For example, if a trader is trading 25% of the total quantity it might be the case that their share of the total cost is simply 25% of the total. We answer this firmly in the negative in Proposition 5.3.
@@ 6 Centralized trading
$\underbrace{\lambda_{1},\dots,\lambda_{n_{1}}}_{\text{Firm target qty: $\lambda_{n_{1}}$}}\quad\text{and}\quad\underbrace{\lambda_{n_{1}+1},\dots,\lambda_{n_{2}}}_{\text{Non-firm tgt qty: $\lambda_{n_{2}}$}}$ (6.1)
Each trader within the firm has a target quantity and associated parent order associated they would like to execute (see Section 2.2. If the firm funnels all the parent orders through an internally operated trading desk, the central trading desk, then it is said to be centralizing its trades. A prominent reason for considering trade centralization is that it might reduce implementation costs. The thinking behind this is that if each trader trades independently then they will operate to minimize their own implementation costs.
@@
Given this setup there are two possible approaches the firm can take. The first case is that they can allow each trader to trade independently (the no-centralization approach) and the second is the firm can centralize all the trades (the cooperative approach). We summarize as in the following table:
@@
No centralization $n$ $n_{1}$ $\lambda_{n_{1}}$ $n_{2}$ $\lambda_{n_{2}}$
Centralization $n_{2}+1$ 1 $\lambda_{n_{1}}$ $n_{2}$ $\lambda_{n_{2}}$
Table 1: Firm and non-firm trading setup as it pertains to how the parameters of competitive trading change when moving from non-centralized to centralized trading.
@@ 6.1 Naive centralization
When a firm naively centralizes it implicitly takes the view that the best way to minimize implementation costs is to trade the centralized order flow as a single monolithic strategy designed to minimize its total implementation cost. In this section, then, we study what this strategy looks like in equilibrium and compare its implementation cost to that of the de-centralized strategies (also in equilibrium). The parameters of the trading game immediately become those of the cooperative row of Table 1. Given this, there are two instances of the firm strategy: one when the traders trade independently and the other where their trades are centralized. We refer to these as the non-centralized and centralized strategies respectively.
@@
$\lambda_{n_{1}}a_{n_{1}}$ $=\lambda_{n_{1}}(B_{n_{1}}(e^{\kappa t}-1)+D_{n_{1}}(1-e^{-\hat{\alpha}t})),\qquad\text{\em Centralized firm strategy}$ (6.2b)
@@
and $\alpha=\kappa\frac{n-1}{n+1}$, $\hat{\alpha}=\kappa\frac{n_{2}}{n_{2}+1}$ and $n_{1}+n_{2}=n$. Next there are the corresponding strategies for the non-firm traders:
@@
In Figure 6 we display comparisons of the firm and non-firm strategies in the centralized versus the non-centralized cases. In each plot green lines represent equilibrium aggregated strategies when firm traders centralize their trades (the cooperative case) and blue lines represent the same but for the non-centralized case. Solid lines represent the firm’s aggregated strategy in equilibrium and dashed lines represent the non-firm traders’ aggregated strategy. We can think of a pair of lines of the same color (one solid and one dashed) as a strategy and a counter-strategy. We now discuss the meaning of this.
@@
Figure 7: An example of firm versus non-firm aggregated strategies in centralized versus non-centralized cases, see Figure 6.
Figures 6 and 7 suggest a counter-intuitive hypothesis. Firms that consolidate trades may not be gaining by doing so. We examine this further next.